OpenStudy (anonymous):

sharon adopted two puppies jack weighs 12 pounds Jill weighs 8 1/3 how heavier is jack than jill plz help me

OpenStudy (anonymous):

and ?

OpenStudy (anonymous):

i think you should make a weight ratio by dividing Jack`s weight over jill`s weight

OpenStudy (anonymous):

is it clear ??

OpenStudy (anonymous):

no you have to show your work

OpenStudy (anonymous):

look how havier jack than jill = \[\frac{ jack`s weight }{ jill`s weight }\]

OpenStudy (anonymous):

=\[\frac{ 12 }{ 8\frac{ 1 }{ 3 } }\]

OpenStudy (anonymous):

how about now ?? its very simple